# How Well Do Large Language Models Serve as End-to-End Secure Code Producers?

### v1

@@ 7. Implications and Discussions
Our study identifies several important implications and suggestions for the research of large language models for code and vulnerability repair.
A. The need for a larger coverage of semantic code analysis engines for vulnerability detection
Semantic code analysis engines such as CodeQL are renowned for their reliability in identifying vulnerabilities. These tools, driven by manually written query scripts targeting specific weaknesses, typically exhibit low false positive rates in practical scenarios. Therefore, they are widely used in research for software security (Pearce et al., 2022a, b). However, current engines fall short in terms of their coverage (i.e., the number of detectable CWEs). Consequently, many studies, including ours, resort to manual code review to ensure comprehensive coverage, albeit at the expense of time and effort. Expanding the coverage of these analysis engines would significantly boost the efficiency and reproducibility of relevant research endeavors.
B. LLMs’ awareness of security risks
In RQ1, it was observed that large language models produced a significant amount of insecure code when tasked with scenarios involving specific security risks. However, the result does not necessarily imply that LLMs are incapable of generating more secure code. One piece of evidence is their ability to correct many of the vulnerabilities present in their generated code when prompted to do so. We posit that the production of vulnerable code by LLMs largely stems from their lack of awareness regarding security issues, as they primarily prioritize fulfilling functional requirements. In real-world scenarios, software developers do not always provide LLMs with information about relevant risks. Therefore, it is crucial to enhance the scenario-relevant security awareness of LLMs, particularly that of code language models which are designed for code-related tasks. Additionally, it is recommended that users explicitly include brief descriptions of potential security weaknesses in prompts to guide LLMs in preventing them.
C. Self-repair ”blind spots” of LLMs
An intriguing observation from RQ3 is that both GPT-3.5 and GPT-4 achieve their lowest success rates when repairing code generated by themselves (as revealed in Table 3). This suggests that similar to human programmers who tend to overlook the weaknesses in their self-written source code, LLMs also exhibit ”blind spots” in code self-repairing. We presume that the phenomenon exists because LLMs are too dependent on the programming patterns learned from their training stage that they tend to ”insist” on these patterns rather than exploring alternative approaches, leaving vulnerabilities unfixed when prompted to address self-produced weak code. Conversely, when fixing insecure code generated by other models, a large language model can leverage its unique patterns to address weaknesses that other LLMs may overlook, resulting in a slightly higher success rate of repair.
D. General large language models versus code language models
In the studies of RQ2, RQ3, and RQ4, we excluded Code Llama and CodeGeeX2 due to their inability to generate responses coherent with our prompts. Although these two language models achieve remarkable results on code generation tasks (hum, 2024), they perform poorly on other code-related tasks such as vulnerability detection and repair, often generating either garbled code or self-conflicting responses. In contrast, general-purpose large language models like GPT-3.5 and GPT-4 can comprehend prompts for detection and repair, thus generating satisfactory results. The disparity may be attributed to the fact that larger-scale language models are trained on extensive natural language datasets, which enables them to comprehend prompts and generate coherent responses. Accordingly, future frameworks for automated secure code construction may either leverage the advantages of general-purpose large language models or utilize specialized models that have been fine-tuned for vulnerability detection and repair.
@@ 8. Threats to Validity
1) Reproducible Code Generation. As generative models, LLMs used in this work are unable to produce completely reproducible output. Given the time-consuming nature of manual code review, we instructed the LLMs to generate only one output for each task (otherwise the amount of code to review would multiply). Consequently, our results may be affected by random factors, as LLMs can produce different outputs when given the same prompt. We contend that such influence is minimal, as all results were generated under default parameters with medium model temperatures. To alleviate doubts, we particularly had GPT-4 generate three parallel outputs for each generation task in RQ1. Manual review confirms that GPT-4 consistently produced similar results across these outputs777The code and review results of this process are included in our public repository.
2) Choice of the Dataset. The SecurityEval dataset used in our work was released two years prior to our research and might have been included in the training data of LLMs. Despite this possibility, all 4 tested LLMs exhibit poor performance in terms of security quality when assessed against this benchmark. Therefore, the security problem of large language models for code remains an open challenge. Furthermore, our conclusions drawn from the experiments with SecurityEval maintain their validity and relevance, as they were not compromised by potential prior exposure of the models to the dataset.
3) Limitations of Code Review. While manual code review is a prevalent practice in the software engineering community, it does not guarantee 100% accuracy because of subjective factors. Moreover, the security of code remains an open question, as standards for secure coding may evolve with the emergence of new attack methods or updates to library functions (e.g., Python standard modules). Nevertheless, we are confident that our method for code review described in 2.3.3 provides largely reliable judgments. Additionally, we make the review results publicly accessible in our repository to mitigate this potential threat.
4) Limitations of Experimental Design. Due to time constraints, we did not fully explore the potential of LLMs in terms of prompt engineering. It is conceivable that prompts with more detailed task descriptions, such as specifying the row and column numbers of vulnerabilities or using different styles, could influence LLM performance. This remains an open question that could be investigated in future research.
@@ 9. Related Work
(1) Vulnerabilities in LLM-generated code. With large amounts of code being generated by LLMs and deployed (sometimes without thorough examination) into production environments every day, their security has become a significant concern for both academia and industry. Pearce et al. evaluated the security of C and Python code generated by GitHub Copilot across 25 CWE cases, finding that 40% of the code was vulnerable (Pearce et al., 2022a). Similarly, Khoury et al. assessed GPT-3.5’s ability to generate code in multiple programming languages for security-critical scenarios and found it failed to meet secure coding standards in 16 out of 21 tasks (Khoury et al., 2023). Additionally, Nair and coauthors demonstrated that ChatGPT produces insecure hardware code if not carefully prompted (Nair et al., 2023). A more recent study by Tihanyi et al. examined the security of C code generated by GEMINI-pro, GPT-4, and other models, revealing that at least 63.47% of the generated programs were vulnerable (Tihanyi et al., 2024), which is a number close to our findings on Python. These results highlight the inability of current large language models to consistently generate secure code without elaborately designed prompts.
@@
(2) LLMs for detecting vulnerabilities in real-world code. Recent research has increasingly focused on the direct application of LLMs in enhancing code security, particularly in the areas of vulnerability detection and repair (Zhou et al., 2024a). Fu et al. investigated the ability of LLMs to detect and classify weaknesses in real-world code (Fu et al., 2023). Purba et al. applied 4 well-known LLMs to detect vulnerabilities in 2 datasets (code gadgets (Li et al., 2018) and CVEfixes (Bhandari et al., 2021), both derived from real-world programs). They found a significant performance gap between the studied LLMs and static analysis tools, primarily due to the high false positive rates of LLMs (Purba et al., 2023), a result consistent with our conclusion in RQ2. Other research also highlighted the limitations of current LLMs in vulnerability detection compared to static analysis tools or specially trained, deep learning-based models (Steenhoek et al., 2024; Yin et al., 2024). Contrary to these findings, some researchers have observed LLMs’ superiority in specific experimental settings. Zhou et al. (Zhou et al., 2024b) and Akuthota et al. (Akuthota et al., 2023) reported better performance of LLMs in certain scenarios. Ullah et al. designed SecLLMHolmes, an automated evaluation framework that investigates LLMs’ reliability in identifying and reasoning about security-related bugs (Ullah et al., 2024). To compensate for the shortcomings of large language models in vulnerability detection, Yang et al. proposed a novel framework that reinforces LLMs with deep learning-based techniques, achieving state-of-the-art performance in vulnerability detection (Yang et al., 2024).
@@
(3) LLMs for repairing vulnerabilities in real-world code. Research results vary regarding the efficacy of large language models in vulnerability repair. A study by Pearce et al. indicates that large language models are promising zero-shot vulnerability fixers (Pearce et al., 2022b). Wu et al. highlighted the advantages of Codex over traditional deep learning-based repair tools for addressing CWE weaknesses in Java code (Wu et al., 2023). Le et al. concluded that ChatGPT provides satisfactory repair results for JavaScript code when detailed descriptions of the vulnerabilities are given (Le et al., 2024). Ahmad et al. found that LLMs such as Codex can effectively fix security bugs in Verilog, a hardware programming language (Ahmad et al., 2024). Conversely, Fu et al. reported that LLMs such as GPT-3.5 underperform compared to traditional models like CodeBERT in vulnerability repair tasks (Fu et al., 2023).
@@
It is noteworthy that most previous research focuses solely on LLMs’ efficacy in detecting and fixing vulnerabilities in real-world, manually written code. While these studies provide valuable insights, they do not fully reveal the potential of LLMs to be end-to-end secure code producers who must detect and repair vulnerabilities in the code they themselves generate. This underscores the need for our research specifically targeting the security of LLM-generated code.
@@ 10. Conclusions and Perspectives
In this paper, we seek an answer to the question of how well large language models serve as end-to-end secure code producers. We first investigate the vulnerabilities present in Python source code generated by GPT-3.5, GPT-4, Code Llama, and CodeGeeX2 on the SecurityEval benchmark. Subsequently, we explore LLMs’ potential to independently enhance the security of the code through code self-review and vulnerable code self-repair. Overall, we manually review 1,452 pieces of code (in RQ1 and RQ3) and automatically examine approximately 4,900 pieces of code (in RQ1, RQ3, and RQ4).
Our study reveals several key findings: (1) large language models tend to generate insecure code in security-critical programming tasks because of their shortage of scenario-relevant awareness of potential risks; (2) large language models such as GPT-3.5 and GPT-4 are not capable of accurately identifying vulnerabilities in the source code they produce, primarily due to their high false positive rates; (3) advance LLMs can achieve up to a 60% success rate repairing insecure code generated by other LLMs, but they exhibit relatively poor performance when repairing self-produced code; (4) Leveraging semantic code analysis engines, a feedback-driven self-iterative repair approach of LLMs significantly enhances the security of LLM-generated code.
While we hold the belief that future large language models have the potential to produce secure code in an end-to-end fashion, current models are unable to accurately fix vulnerable code without assistance from established tools like semantic code analysis engines.
@@
Appendix A APPENDIX
A.1. Detailed results of RQ2
Figure 14 and Figure 15 respectively show the detailed results of GPT-3.5 and GPT-4 in detecting vulnerabilities in code generated by the four studied large language models. The results in the colored cells are presented as for each CWE. For example, the first cell of Figure 15 indicates that GPT-4 correctly identified whether the code is vulnerable to CWE-20 (Improper Input Validation) for 3 out of 6 pieces of code generated by GPT-3.5.
A.2. Detailed results of RQ3
Figure 16 and Figure 17 respectively show the detailed results of GPT-3.5 and GPT-4 in fixing vulnerabilities in code generated by the four studied large language models. The results in the colored cells are presented as for each CWE. For example, the first cell of Figure 17 indicates that GPT-4 successfully fixed all the 3 vulnerable pieces of code generated by GPT-3.5 with the vulnerability of CWE-20. A gray cell indicates that no code with that CWE weakness was generated in the first place.

### v2

@@ 7. IMPLICATIONS AND DISCUSSIONS
7.1. LLMs’ awareness of security risks
In RQ1, it was observed that LLMs produced a significant amount of insecure code when tasked with scenarios involving specific security risks. However, the result does not necessarily imply that LLMs are incapable of generating more secure code, for they were able to correct many of the vulnerabilities present in their generated code when prompted to do so. We posit that the production of vulnerable code by LLMs largely stems from their lack of awareness regarding security issues, as they primarily prioritize fulfilling functional requirements. In real-world scenarios, software developers do not always provide LLMs with information about relevant risks. Therefore, it is crucial to enhance the scenario-relevant security awareness of LLMs, particularly that of code language models which are designed for code-related tasks. Additionally, it is recommended that users explicitly include brief descriptions of potential security weaknesses in prompts to guide LLMs in preventing them.
7.2. Self-repair “blind spots” of LLMs
An intriguing observation from RQ3 is that both GPT-3.5 and GPT-4 achieve their lowest success rates when repairing code generated by themselves (as revealed in Table 4). This suggests that similar to human programmers who tend to overlook the weaknesses in their self-written source code, LLMs also exhibit “blind spots” in code self-repairing. We presume that the phenomenon exists because LLMs are too dependent on the programming patterns learned from their training stage that they tend to “insist” on these patterns rather than exploring alternative approaches, leaving vulnerabilities unfixed when prompted to address self-produced weak code. Conversely, when fixing insecure code generated by other models, a large language model can leverage its unique patterns to address weaknesses that other LLMs may overlook, resulting in a slightly higher success rate of repair.
@@ 8. THREATS TO VALIDITY
Choice of the Dataset: The SecurityEval dataset used in our work was released two years prior to our research and might have been included in the training data of LLMs. Despite this possibility, all 4 tested LLMs exhibit poor performance in terms of security quality when assessed against this benchmark. Therefore, the security problem of large language models for code remains an open challenge and our conclusions drawn from the experiments with SecurityEval maintain their validity and relevance.
Limitations of Code Review: While manual code review is a prevalent practice in the software engineering community, it does not guarantee 100% accuracy because of subjective factors. Moreover, the security of code remains an open question, as standards for secure coding may evolve with the emergence of new attack methods or updates to library functions (e.g., Python standard modules). Nevertheless, we are confident that our method for code review described in 2.3.3 provides largely reliable judgments. Additionally, we make the review results publicly accessible in our repository.
Limitations of Experimental Design: Due to time constraints, we did not fully explore the potential of LLMs in terms of prompt engineering. It is conceivable that prompts with more detailed task descriptions, such as specifying the row and column numbers of vulnerabilities or using different prompting styles, could influence the performance of LLMs.
@@ 9. RELATED WORK
(1) Vulnerabilities in LLM-generated code. Pearce et al. were among the first to evaluated the security of C and Python code generated by GitHub Copilot across 25 CWE cases, finding that 40% of the code was vulnerable (Pearce et al., 2022). Furthermore, (Khoury et al., 2023; Asare et al., 2023; Sandoval et al., 2023; Fu et al., 2023a; Liu et al., 2024) also examine LLMs’ ablity to generate secure code in various settings. A more recent study by Tihanyi et al. examined the security of C code generated by GEMINI-pro, GPT-4, and other models, revealing that at least 63.47% of the generated programs were vulnerable (Tihanyi et al., 2024), which is a number close to our findings on Python. These results highlight the inability of current LLMs to consistently generate secure code without elaborately designed prompts.
@@
(2) LLMs for detecting vulnerabilities in real-world code. Recent research has increasingly focused on the direct application of LLMs in enhancing code security, particularly in the areas of vulnerability detection and repair (Zhou et al., 2024a). Fu et al. investigated the ability of LLMs to detect and classify weaknesses in real-world code (Fu et al., 2023b). Purba et al. applied 4 well-known LLMs to detect vulnerabilities in 2 datasets. They found a significant performance gap between the studied LLMs and static analysis tools, primarily due to LLMs’ high false positive rates (Purba et al., 2023), a result consistent with our conclusion in RQ2. Other research also highlighted the limitations of current LLMs in vulnerability detection compared to static analysis tools or specially trained, deep learning-based models (Steenhoek et al., 2024; Yin et al., 2024). Contrary to these findings, some researchers have observed LLMs’ superiority in specific experimental settings (Zhou et al., 2024b)(Akuthota et al., 2023). Ullah et al. designed SecLLMHolmes, an automated evaluation framework that investigates LLMs’ reliability in identifying and reasoning about security-related bugs (Ullah et al., 2024). Yang et al. proposed a novel framework that reinforces LLMs with deep learning-based techniques, achieving state-of-the-art performance in vulnerability detection (Yang et al., 2024).
@@
(3) LLMs for repairing vulnerabilities in real-world code. Research results vary regarding the efficacy of large language models in vulnerability repair. A study by Pearce et al. indicates that LLMs are promising zero-shot vulnerability fixers (Pearce et al., 2023). Wu et al. highlighted the advantages of Codex over traditional deep learning-based repair tools for addressing CWE weaknesses in Java code (Wu et al., 2023). Le et al. concluded that ChatGPT provides satisfactory repair results for JavaScript code when detailed descriptions of the vulnerabilities are given (Le et al., 2024). Ahmad et al. found that LLMs such as Codex can effectively fix security bugs in Verilog, a hardware programming language (Ahmad et al., 2024). Conversely, Fu et al. reported that LLMs such as GPT-3.5 underperform compared to traditional models like CodeBERT in vulnerability repair tasks (Fu et al., 2023b).
@@
It is noteworthy that most previous research focuses solely on LLMs’ efficacy in detecting and fixing vulnerabilities in real-world, manually written code. While these studies provide valuable insights, they do not fully reveal the potential of LLMs to be end-to-end secure code agents who must detect and repair vulnerabilities in the code snippets generated by themselves. This underscores the need for our research specifically targeting the security of LLM-generated code.
@@ 10. CONCLUSIONS AND PERSPECTIVES
In this paper, we seek an answer to the question of how well LLMs serve as end-to-end secure code agents for Python. We first investigate the vulnerabilities in Python source code generated by GPT-3.5, GPT-4, Code Llama, and CodeGeeX2 on the SecurityEval benchmark. Subsequently, we explore LLMs’ potential to independently enhance the security of the code through code self-review and vulnerable code self-repair. Overall, we manually or automatically examined around 4,900 code pieces.
Our study reveals several key findings: (1) LLMs tend to generate insecure code in security-critical programming tasks because of shortage of scenario-relevant awareness of potential risks; (2) LLMs such as GPT-3.5 and GPT-4 are not capable of accurately identifying vulnerabilities in the source code they produce, primarily due to their high false positive rates; (3) advance LLMs can achieve up to a 60% success rate repairing insecure code generated by other LLMs, but they exhibit relatively poor performance when repairing self-produced code; (4) Leveraging semantic code analysis engines, a feedback-driven self-iterative repair approach significantly enhances the security of LLM-generated code.
While we hold the belief that future LLMs have the potential to produce secure code in an end-to-end fashion, current models are unable to serve as reliable secure code agents without assistance from established tools like semantic code analysis engines.
@@
11. ACKNOWLEDGMENTS
In the preparation of this manuscript, ChatGPT was utilized to refine certain sentences to enhance their clarity, precision, and academic rigor. However, the core ideas, analysis, and conclusions remain entirely the authors’ own. The use of AI-assisted editing was strictly limited to linguistic improvements and did not influence the scholarly integrity or originality of the research.